A gardener is making a triangular planting, with 46 plants in the front row, 41 in the second row, 36 in the third row, and so on. If the pattern is consistent, how many plants are there altogether?
Answer by richwmiller(17219)

an = a + (n - 1)d
1 = 41 + (n - 1)-5
-40/-5=n-1
8=n-1
n=9
S = ½(a + an)n
S = ½(41 + 1)9
S=1/2(42)*9
s=21*9
s=189 plants altogether